### v1

# Jump-teaching: Ultra Efficient and Robust Learning with Noisy Label

Abstract
Sample selection is the most straightforward technique to combat label noise, aiming to distinguish mislabeled samples during training and avoid the degradation of the robustness of the model. In the workflow, selecting possibly clean data and model update are iterative. However, their interplay and intrinsic characteristics hinder the robustness and efficiency of learning with noisy labels: 1) The model chooses clean data with selection bias, leading to the accumulated error in the model update. 2) Most selection strategies leverage partner networks or supplementary information to mitigate label corruption, albeit with increased computation resources and lower throughput speed. Therefore, we employ only one network with the jump manner update to decouple the interplay and mine more semantic information from the loss for a more precise selection. Specifically, the selection of clean data for each model update is based on one of the prior models, excluding the last iteration. The strategy of model update exhibits a jump behavior in the form. Moreover, we map the outputs of the network and labels into the same semantic feature space, respectively. In this space, a detailed and simple loss distribution is generated to distinguish clean samples more effectively. Our proposed approach achieves almost up to 2.53×2.53$\times$2.53 × speedup, 0.46×0.46$\times$0.46 × peak memory footprint, and superior robustness over state-of-the-art works with various noise settings.
11footnotetext: Equal contribution22footnotetext: Correspondence to chengfei@xidian.edu.cn
1 Introduction
Learning with Noisy Label (LNL) is the most promising technique in weakly supervised learning. Generally, noisy labels stem from mistaken annotations of the dataset, such as in crowd-sourcing[39] and online query[4] As accurate annotations of large datasets are a time-consuming endeavor, the existence of noisy labels becomes inevitable. Deep neural networks can easily overfit to noisy labels, which is prone to poor generalization performance[49, 13]. Furthermore, the efficiency challenge of LNL is often overlooked in comparison to the robustness problem[3], which is vital in real-time[27, 3] or security scenarios[1, 52, 9]. Thus, advanced methods for LNL are essential.
Recent approaches of LNL can be classified into three types: regularization, label correction, and sample selection. Regularization methods focus on crafting noise-robust loss functions[36, 10, 37, 50] and regularization techniques[23, 51, 5]. These methods cannot fully distinguish noisy samples during training, resulting in suboptimal outcomes. Label correction techniques, integrating closely with semi-supervised learning, aim to refine or recreate pseudo-labels [35, 14, 33, 30]. These methods effectively use noisy samples but typically require substantial computational resources.
(a) Sym. $\epsilon=0.5$
(b) Sym. $\epsilon=0.8$
(c) Sym. $\epsilon=0.5$
(d) Sym. $\epsilon=0.8$
Figure 1: Test accuracy and "disagreement" of different strategies with symmetric noise ratio $\epsilon=0.5$, $\epsilon=0.8$. We use Intersection over Union (IoU) between two selections to represent the disagreement. In the cross-update strategy, the disagreement emerges in a different network, while in the other two strategies, it emerges during different epochs. More details are discussed in Appendix A.11
Sample selection can be regarded as a potential method for LNL[34, 19, 40, 28, 12, 42]. For a given task, it tends to first select possibly clean samples by a certain strategy and trains on them with iterations. When this strategy is appropriately configured, more clean samples are acquired during training, resulting in a more robust neural network. Unfortunately, selection strategies are always accompanied by sample-selection bias[12]. This bias introduces noisy samples to each training data. When the neural network trains on these data, the error erode its robustness. Consequently, as the training iterations increase, the rate of error accumulation accelerates. In other words, the error flows from the beginning to the end of training through iterations. In order to reduce the error accumulation, many works are proposed, such as Decoupling[28], Co-teaching[12] and Co-teaching+[47], etc. These methods follow a similar paradigm: an extra network is integrated to give different predicted labels from the original network. These disagreements between the extra and original network try to diverge in each sample selection and guide the training process by model updating.
However, these methodologies fail to strike a suitable equilibrium between efficiency and robustness, making them unsuitable for real-time and security-intensive situations. These challenges have long been ignored for a variety of reasons. Firstly, existing works are designed with consideration only for robustness. However, the extra network commonly used in such methods requires significant computation resources, and thus neglects the efficiency. Intuitively, if we can find the "disagreement" in a single neural network, without any extra network, and exploit it, it is possible to find a balance between efficiency and accuracy. Secondly, the oversimplified design of the loss function leads to the ineffective selection of clean samples. Generally, the loss function calculates the discrepancy between the predicted probabilities and the one-hot encoding, which outputs a floating-point value. This design affects the results of sample selection and learning tasks in LNL. In addition, a single floating-point value provides insufficient information for effective sample selection.
In this paper, we propose an ultra-efficient and robust learning method with noisy labels called "Jump-teaching", which bridges the disagreement between different training iterations in a single network and enriches the loss representation of sample selection with an auxiliary and lightweight head at the last layer of the network. Concretely, we first discover that there is a disagreement between different training iterations of a neural network, excluding neighboring iterations, from a temporal perspective in Fig. 1. Therefore, This disagreement enables the prevention of error accumulation caused by selection bias. Meanwhile, we design an auxiliary and lightweight head at the last layer of the network only for sample selection. This head transforms the predicted probability and the one-hot encoding into semantic features and hash coding, respectively, in other spaces.When training on clean data, the goal of this head is to minimize the gap between semantic features and hash coding as close to 0 (a zero vector) as possible. The ideal distribution of elements of this vector is a mean value distribution with a variance of 00. Conversely, for noisy data, this gap becomes too large to reach $\mathbf{0}$, and its distribution is uncertain. According to the aforementioned characteristics, we can easily and effectively distinguish between clean and noisy samples by this particular head.
Our contributions can be summarized as follows: 1) We exploit the disagreement between intervalic iterations during training for a single network, and leverage the jump-update strategy to suppress the accumulated error. To the best of our knowledge, it is the first time that a single network has surpassed the capacity of a dual network in the robustness and efficiency of LNL. (2) The loss of sample selection is proposed with a lightweight design and easily selects clean data by intrinsic characteristics. 3) Our proposed approach achieves almost up to 2.53×2.53$\times$2.53 × speedup, 0.46×0.46$\times$0.46 × peak memory footprint, and outperforms state-of-the-art methods in terms of robustness with various noise settings, especially for real-time scenarios. Besides, there is considerable flexibility in integrating this method with other LNL approaches.
The rest of this paper is organized as follows. Related work is reviewed in Sec. 2. In Sec. 3, we introduce our ultra-efficient and robust algorithm Jump-teaching. Experiments are illustrated in Sec. 4. Conclusions are given in Sec. 5.
2 Related Work
Sample Selection. The core idea of the sample selection approach is to filter out noisy samples to prevent the network from fitting to them. Since the loss of a single sample is insufficient for sample selection, current approaches require additional information. Most works employ a dynamic threshold with foreknowledge of the noise ratio, e.g., Co-teaching[12] and Co-teacing+[47], or probabilistic estimation on loss value, such as Gaussian Mixture Models (GMM) [29] and Beta Mixture Models [26]. The former always requires prior knowledge, such as noise ratio[12, 47], which doesn’t work in the real world. Other approaches rely on a complex statistical estimation process[21, 38, 2].
Sample-selection Bias. Sample selection inherently involves bias, leading to accumulation error. To avoid this, many methods employ dual networks and correct this bias through the divergence in their selection. Decoupling[28] utilizes a teacher model to select clean samples to guide the learning of a student model, Co-teaching[12] simultaneously trains two networks on data selected by peer network. Co-teaching+ [47] maintains divergence between the two networks by limiting the training data to samples where the networks disagree. JoCoR[38] trains two networks with the same data and employs the regularization technique to remain divergent. Additionally, many LNL methods integrate the co-training framework to achieve state-of-the-art performance[21, 34, 23, 24, 6]. However, these co-training methods significantly increase computational and storage overhead.
3 Methodology
Jump-teaching is an ultra-efficient and robust sample selection method with only one network for noisy labels. This algorithm employs Jump-update Strategy to update the model and Semantic Loss Decomposition to select clean data. We demonstrate them in Sec. 3.1 and Sec. 3.2, respectively.
3.1 Jump-update Strategy
(a) Model Udpates Strategy
(b) Error Flow
Figure 2: As errors accumulate throughout the process, many solutions are suggested to minimize these errors with different types of error flow.
In this section, we provide a robust and efficient update paradigm known as the jump-update strategy. Furthermore, we offer both theoretical and experimental analysis to explain its effectiveness and then explore its relationship with other approaches.
Strategy Description sTo simplify our discussion, we dive into the procedure of sample selection and give some definitions. Sample selection is accomplished by the execution of model updates and clean sample selection with multiple iterations. In other words, model updates and clean sample selection are performed only once in every iteration. Previous strategies and ours are shown in Fig. 2(2(a)). Suppose $t_{i}$ denotes the current $i$-th iteration, ancestor iteration refers to the former iteration $t_{j},0\leq j\leq{i-2}$, excluding the previous $(i-1)$-th iteration, during the procedure of sample selection. Similarly, descendant iteration denotes the future iteration $t_{s},i+1\leq s\leq N_{iterations}$. $N_{iterations}$ is the total number of training iterations. In our paradigm, the current network in the $i$-th iteration is trained with clean samples selected only by the network from an ancestor iteration $t_{j}$. This behavior of sample selection exhibits a jump form. The origin of the algorithm’s name is derived from this point. Concretely, we leverage a binary identifier to represent the outcome of the label judgment after clean sample selection. Thus, a binary identification table corresponds to the entire data. The jump-update strategy is divided into three steps: 1) The neural network in current interaction $t_{i}$ generates the new binary identification table by the clean sample selection for the descendant iteration $t_{s}$. 2) We fetch and cache the old binary identification table from the ancestor iteration $t_{j}$. 3) The network updates the parameters based on the clean data judged only by the old table. Before the update, the network inherits the weight from the previous iteration $t_{i-1}$. When the jump-update strategy is applied, disagreements between different training iterations of a neural network appear, leading to a decrease in the accumulated error. Suppose $T$ denotes the jump steps, it should be noted that $T=i-j=s-i,2\leq T\leq N$. The evidence of disagreement and the setting of $T$ are illustrated in Sec. 4.2.
Theoretical Analysis Sample selection inevitably has a selection bias, leading to accumulated error in an error flow. As shown in Fig. 2(2(b)), the graph of error flow presents a sequential form by iterations while the jump-update strategy splits a sequential error flow into multiple error sub-flows. Intuitively, the more training iterations increase, the more rapidly the error is accumulated. Nonetheless, the degree of error accumulation in the two is significantly different. In the sequential form, errors are accumulated consecutively. As the error sub-flows are orthogonal in the jump-update strategy, each error is accumulated only in its own error sub-flow. Thus, the jump-update strategy has a significantly smaller degree of accumulated error compared to the sequential form. This is the source of the magic of the jump-update strategy. We formalize the aforementioned procedure mathematically below and give some detail properties.
Suppose that $N_{A}$ is the total number of error accumulations, $N_{a}$ is the number of error accumulations in an error subflow, and $N_{f}$ is the number of error sub-flows. Besides, the constant $e$ represents the number of training epochs and $n$ represents how many selections are made in one epoch. $D_{A}$ denotes the overall degree of accumulated error, while $d^{k}_{a}$ denotes the degree of accumulated error in the $k$-th error sub-flow. In the absence of a specific reference, we can also denote the degree of accumulated error in one error sub-flow by $d_{a}$.
Property 1
The overall degree of accumulated error $D_{A}$ is proportional to the total number of error accumulation $N_{A}$, $D_{A}\propto N_{A}$. The number of error accumulation $N_{A}$ equals the total number of training iterations $N_{iterations}$, while $N_{iterations}$ equals $e\times n$. Therefore, $D_{A}\propto n$.
The overall degree of accumulated error $D_{A}$ depends on $n$ stated in Property 1. When the network selects data from a mini-batch, $D_{A}$ can be enormous because $n$ is equal to the number of mini-batches, e.g., Co-teaching. If the network selects data from the entire dataset, $n$ can be reduced to $\mathbf{1}$, thereby significantly reducing $D_{A}$, e.g., TopoFilter [40]. However, this is inefficient because an additional forward over the entire dataset before training is necessitated. Property 1 is proven in Appendix A.1.
Property 2
The accumulated error could be reduced by splitting the error flow into multiple error sub-flows. The degree of accumulated error in the $k$-th error sub-flow $d^{k}_{a}$ is proportional to the number of error accumulations in each error sub-flow $N_{a}$, $d^{k}_{a}\propto N_{a}$. The number of error accumulations in each error flow $N_{a}$ equals to $\frac{N_{A}}{N_{f}}$.
Property 1 states that $D_{A}$ can be reduced by minimizing $n$, while $D_{A}$ can also be reduced by splitting into error sub-flows is illustrated in Property 2. The first property relates to the sample selection mechanism, and the second property is associated with the different strategies of model updates: 1) The self-update strategy follows a single error flow, resulting in $d_{a}=D_{A}$, which leads to rapid error accumulation. 2) The cross-update strategy has two error sub-flows and $d_{a}=0.5D_{A}$, which mitigates the error accumulation to some extent. 3) The jump-update strategy reduces $d_{a}$ to $0.5e$, which is a significantly minimal value. Further insights are detailed in Appendix A.2.
Experimental Analysis
(a) $D_{A}$ with different $n$
(b) $D_{A}$ with different strategies
[…]
We verify Property 1 and Property 2 with toy examples, respectively. We employ the small-loss sample selection method from Co-teaching to establish the baselines for self-update and cross-update. Our experiments choose CIFAR-10[20] dataset with the symmetric noise ratio $\epsilon=80\%$. We utilize ResNet-18[16] as the backbone and warm up it for one epoch before formal training.
To verify Property 1, we observe $D_{A}$ by testing the accuracy of the network with different values of $n$. To control $n$, We set $r$ as the proportion of sample selection that takes effects, $D_{A}$ equals $r\times n$ in this way. We set $r$ to $10\%$, $30\%$, $50\%$, $80\%$, and $100\%$. As shown in Fig. 3(3(a)), rapidly accumulated errors lead to extreme deterioration of the model such as when $r=100\%$ and $r=80\%$, while slower accumulated ones achieve better performance, with a moderate $r=50\%$ yielding the best results. This is consistent with Property 1.
To verify Property 2, we observe $D_{A}$ by testing the accuracy of the neural network with three strategies, self-update, cross-update, and jump-update. As shown in Fig. 3(3(b)), cross-update slightly outperforms self-update, while jump-update is significantly more effective than both. Thus, splitting error flows is an effective way to reduce error accumulation. Moreover, we also greatly improved the performance of self-update by reducing $n$. Specifically, we employ two unbiased initial identifier tables, with $r$ set to $30\%$ and $5\%$, respectively. This is discussed in Appendix A.3, which also includes the impact of initial bias.
Conclusion The jump-update strategy is more effective and efficient than previous works. Compared to methods that implicitly leverage Property 1, it can reduce $D_{a}$ to a small number i.e., $0.5e$ at a constant cost. Compared to the cross-update strategy that leverage Property 2, it can not only reduce the degree of accumulated error significantly more but also halve the training cost.
3.2 Semantic Loss Decomposition
Motivation According to memorization effects, neural networks prioritize learning simpler patterns from data [49]. The small-loss criterion leverages this effect: samples with clean labels are learned first by the network, hence exhibiting smaller losses compared to noisy samples. However, the relative magnitude of losses is determined by comparison with other samples e.g., rank sample losses with Top-k algorithm[12, 17] or modeling loss distribution[21, 29]. To avoid such costly overheads, we come up with an idea: Is it possible to decompose the single loss and build a distribution? Inspired by memorization effects, if the flipped labels share some identical components with original labels, then for flipped labels, the clean components will be learned first and incur smaller losses, while the noisy components will result in larger losses. This property can be utilized to identify whether a label contains noise. Therefore, we design an auxiliary and lightweight head at the last layer of the network only for sample selection. This head transforms the predicted probability and the one-hot encoding into semantic feature embedding and hash coding, respectively, in other spaces. This enriches the representation of the loss function for sample selection.
We formally hypothesize the existence of a non-orthogonal codebook capable of transforming one-hot encodings into hash codes. Correspondingly, the network output is mapped to feature embedding. During training, the distance $\mathbf{d}(\mathbf{z}_{t},\mathbf{y})$ between the output $\mathbf{z}_{t}$ of a clean sample at time $t$ and its label $\mathbf{y}$ can be represented as $\lim_{t\to\infty}\mathbf{d}(\mathbf{z}_{t},\mathbf{y})=\mathbf{0}$. The notation $\mathbf{0}$ denotes a vector of zeros, indicating that the distances across different bits uniformly converge towards zero as $t$ approaches infinity.
Hash Coding with Hardamard Matrix Inspired by [45], we utilize the excellent properties of Hadamard matrices to construct mappings for category encoding. A $K$-bit Hadamard matrix can generate $2K$ codewords, each $K$ bits long, with a minimum Hamming distance of $\frac{K}{2}$. For $K$-bit hash codes, we construct a $K\times K$ Hadamard matrix. From this, we select $c$ row vectors as category encodings, each with a Hamming distance of $\frac{K}{2}$. Labels are mapped to the semantic space through this codebook. This mapping can be presented as $H:\tilde{y}_{i}\in\{0,1\}^{C}\rightarrow\tilde{y}_{i}^{\prime}\in\{-1,1\}^{k}$.
Feature Embedding and Loss Function outputs are mapped to feature embeddings in a subspace via an additional three-layer MLP as detection head, which represented as $f:x_{i}\in\mathbb{R}^{n}\rightarrow z_{i}\in\mathbb{R}^{k}$.
In this framework, we employ binary cross-entropy[31] to define the distance $\mathbf{d}$ between the network’s predictions and the labels, which indicates whether the network adequately captures the semantics of each component, which is formulated as
$\mathbf{d}(\mathbf{z}_{i},\tilde{\mathbf{y}}_{i}^{\prime})=-\left[\tilde{\mathbf{y}}_{i}^{\prime}\odot\log(\mathbf{z}_{i})+(1-\tilde{\mathbf{y}}_{i}^{\prime})\odot\log(1-\mathbf{z}_{i})\right]$ (1)
We apply the arithmetic mean for supervising this head, where the loss function can be articulated as
$\mathcal{L}_{i}^{BCE}=-\frac{1}{K}\sum_{j=1}^{K}\left[\tilde{y}_{ij}^{\prime}\log(z_{ij})+(1-\tilde{y}_{ij}^{\prime})\log(1-z_{ij})\right]$ (2)
The distance vector $\mathbf{d}(\mathbf{z}_{i},\tilde{\mathbf{y}}_{i}^{\prime})$ obtained through semantic decomposition directly describes the distribution of sample loss in semantic space. Compared to a single loss value, it provides richer information. To leverage the network’s memorization effect, we use variance to characterize the disparity in the network’s learning degree across different semantic components.
$\mathrm{Var}(\mathbf{d}(\mathbf{z}_{i},\tilde{\mathbf{y}}_{i}^{\prime}))=\frac{1}{K}\left(\mathbf{d}(\mathbf{z}_{i},\tilde{\mathbf{y}}_{i}^{\prime})-\mathcal{L}_{i}^{BCE}\mathbf{1}\right)^{T}\left(\mathbf{d}(\mathbf{z}_{i},\tilde{\mathbf{y}}_{i}^{\prime})-\mathcal{L}_{i}^{BCE}\mathbf{1}\right)$ (3)
We set a fixed threshold to distinguish clean samples from noisy samples, which is independent of other samples and different training phases. The identifier of a sample is updated as follows:
$I_{i}=\begin{cases}\text{True}&\text{if }\mathrm{Var}(\mathbf{d}(\mathbf{z}_{i},\tilde{\mathbf{y}}_{i}^{\prime}))\leq\tau,\\ \text{False}&\text{otherwise}.\end{cases}$ (4)
Since this threshold is expected to approach zero infinitely, we set it to 0.001.
3.3 Training Pipeline
Algorithm 1 Jump-teaching
0:  network parameters $w_{f}$, learning rate $\eta$, maximum epochs $T_{\text{max}}$, maximum samples $S_{\text{max}}$, identifier table $I$
0:  $w_{f}$
1:  Initialize $I[i]\leftarrow\text{True}$ for all $i$
2:  for $t=1$ to $T_{\text{max}}$ do
3:     Permute $D$ randomly
4:     for $s=1$ to $I_{\text{max}}$ do
5:        $clean\_flag\leftarrow I[s]$
6:        Fetch current sample $D[s]$
7:        Update $I[s]$ based on Eq. 7
8:        if $clean\_flag$ then
9:           Update: $w_{f}\leftarrow w_{f}-\eta\nabla\mathcal{L}^{BCE}(w_{f},D[s])$
10:           Update: $w_{f}\leftarrow w_{f}-\eta\nabla\mathcal{L}^{CE}(w_{f},D[s])$
11:        end if
12:     end for
13:  end for
14:  return  $w_{f}$
Semantic Loss Decomposition should be considered an effective method for finding clean samples. In this section, we discuss how it can assist in selection. During training, a three-layer MLP is employed as an auxiliary noise detection head on top of the existing network. It is trained by Eq. 2. Meanwhile, the network’s classification head continues to train normally and is used for inference.
However, the different convergence rates of the two learning heads impede the model’s robust learning capabilities. The cross-entropy loss, the objective function for classification tasks, is more readily optimizable and thus converges significantly faster than the detection head and can lead to premature over-fitting of noise, resulting in error accumulation. In order to balance the convergence rate of the two, we employ temperature scaling to calibrate the label probabilities[11] of the classification head. Thus, the soften softmax function will be:
$\sigma_{\mathrm{softmax}}\left(\mathbf{x}_{ij}\right)=\frac{\exp\left(\mathbf{x}_{ij}/T\right)}{\sum_{j=1}\exp\left(\mathbf{x}_{ij}/T\right)}$ (5)
Inspired by [43], we also make use of a classifier head to make simple selection based on that the predictions of the model are consistent with the labels. The new clean table $I$ is then evaluated as:
$I=(\hat{y}_{i}==\tilde{y}_{i})$ (6)
where $\hat{y}_{i}=\arg\max_{j}p_{i}^{j}$ is the prediction label and $p_{i}^{j}$ represents the probability of the $j$-th class for the $i$-th sample. $\tilde{y}_{i}$ denotes the label of the $i$-th sample. Finally, we can update the table by combine Eq. 4 and Eq. 6:
$I^{\prime}=I_{detection}\lor I_{classifier}$ (7)
The algorithm of Jump-teaching is shown in Algorithm 1.
4 Experiments
In this section, we demonstrate the effectiveness of our proposed approach, Jump-teaching, compared with the state-of-the-art in Sec. 4.1. Ablation study is illustrated in Appendix A.10. As the jump-update strategy and semantic loss decomposition are the two integral components of this approach, we thoroughly examine each of them in Sec. 4.2 and Appendix A.13.
Noisy Benchmark Datasets. We verify the experiments on three benchmark datasets, including CIFAR-10, CIFAR-100 and Clothing1M[44]. These datasets are summarized in Appendix A.4. These datasets are popular for evaluating noisy labels. Following the setup on [21, 23], we simulate two types of label noise: symmetric noise, where a certain proportion of labels are uniformly flipped across all classes, and asymmetric noise, where labels are flipped to specific classes, e.g., $bird\rightarrow{airplane}$, $cat\leftrightarrow{dog}$. Assume $\epsilon$ denotes the noise ratio, their mathematical definitions are in Appendix A.6.
Baselines. To be more convincing, we compare the competing methods of LNL. These methods are as follows: Standard, which is simply the standard deep network trained on noisy datasets, Decoupling[28], Co-teaching[12], Co-teaching+[47], PENCIL[46], TopoFilter[40], ELR[23];, FINE[19], SPRL[32] APL[25], CDR[41], MentorNet[17], SIGUA[13], JoCor[38] and CoDis[42]. A brief overview of available source code of these methods is illustrated in Appendix A.7.
Experimental Settings. All experiments operate on a server equipped with an NVIDIA A800 GPU and PyTorch platform. In the following experiments, Jump-teaching almost employs the same configuration. It trains the network for $200$ epochs by SGD with a momentum of $0.9$, a weight decay of $31e-$ 3, and a batch size of $128$. The initial learning rate is set to $0.2$, and a cosine annealing scheduler finally decreases the rate to $45e-$ 4. The warm-up strategy is utilized by Jump-teaching, and the warm-up period is $30$ epochs. After the warm-up period, we augmented the data as detailed in the Appendix A.5. The threshold of variance $\tau=0.001$, discussed in Appendix A.9.Jump step is stated in Appendix A.12. Exceptionally, we set the weight decay as $45e-$ 4 to facilitate learning on fewer available samples when the noise ratio $\epsilon$ equals $50\%$ and $80\%$ in CIFAR-100, respectively. The single network of Jump-teaching employs three types of backbone networks to fulfill different requirements of the experimental design, such as PreActResNet-18[15], ResNet-18, and three layers of neural network. Moreover, the architectures of these backbones and the auxiliary head are illustrated in Appendix A.8. The baseline methods fully follow the experimental setup in the literature [12, 21].
4.1 Comparison with the State-of-the-Arts
Synthetic Noisy Benchmark. We compare our proposed method with the following representative approaches: Standard, Decoupling, Co-teaching, Co-teaching+, PENCIL, TopoFilter, ELR, FINE, and SPRL. As shown in Table 1, Jump-teaching demonstrates superior performance with different noise settings. With the symmetric noise ratio $\epsilon=0.8$, its accuracy has improved by $13.3\%$ and $16.2\%$ on two datasets. This indicates its strong robustness.
Table 1: Test accuracy(%) on CIFAR-10 and CIFAR-100 with symmetric and asymmetric noise. All methods employ PreActResNet-18 and train $200$ epochs with three trials. The best results are highlighted in bold.
Dataset CIFAR-10 CIFAR-100
Noise Type Sym. Asym. Sym. Asym.
Noise ratio 0.2 0.5 0.8 0.4 0.2 0.5 0.8 0.4
Standard 84.6 $\pm$ 0.1 62.4 $\pm$ 0.3 27.3$\pm$0.3 75.92 $\pm$0.4 56.1 $\pm$0.1 33.6 $\pm$ 0.2 8.2 $\pm$0.1 40.1 $\pm$0.2
Decoupling(’17) 86.4 $\pm$ 0.1 72.9 $\pm$ 0.2 48.4 $\pm$ 0.6 83.3 $\pm$ 0.2 53.3 $\pm$ 0.1 28.0 $\pm$ 0.1 7.9 $\pm$ 0.1 39.9 $\pm$ 0.4
Co-teaching(’18) 89.9 $\pm$ 0.6 67.3 $\pm$4.2 28.1$\pm$2.0 79.2$\pm$0.5 61.8 $\pm$0.4 34.7$\pm$0.5 7.5$\pm$0.5 40.0$\pm$1.2
Co-teaching+(’19) 88.1 $\pm$ 0.0 61.8 $\pm$ 0.2 22.3 $\pm$ 0.6 58.2$\pm$ 0.2 54.5 $\pm$ 0.1 27.6 $\pm$ 0.1 8.4 $\pm$ 0.1 19.87$\pm$0.3
PENCIL(’19) 88.2 $\pm$ 0.6 73.4 $\pm$ 1.5 36.0 $\pm$ 0.7 77.3 $\pm$ 3.4 57.4 $\pm$ 1.0 11.4 $\pm$ 3.2 5.4 $\pm$ 1.2 45.7 $\pm$ 0.9
Topofilter(’20) 89.5 $\pm$ 0.1 84.6 $\pm$ 0.2 45.9 $\pm$ 2.6 89.9 $\pm$ 0.1 63.9 $\pm$ 0.8 51.9 $\pm$ 1.7 16.9 $\pm$ 0.3 66.6 $\pm$ 0.7
FINE(’21) 90.2 $\pm$ 0.1 85.8 $\pm$ 0.8 70.8 $\pm$ 1.8 87.8 $\pm$ 0.1 70.1 $\pm$ 0.3 57.9 $\pm$ 1.2 22.2 $\pm$ 0.7 53.5 $\pm$ 0.8
SPRL(’23) 91.7 $\pm$ 0.2 88.35 $\pm$ 0.6 63.9 $\pm$ 1.4 89.8 $\pm$ 0.5 69.5 $\pm$ 0.8 57.17 $\pm$ 1.6 23.8 $\pm$ 2.2 58.72 $\pm$ 1.3
Jump-teaching 94.8$\pm$0.1 92.2$\pm$0.1 84.1 $\pm$ 1.1 90.7 $\pm$0.3 72.7$\pm$ 0.5 67.1$\pm$0.2 40.0 $\pm$1.1 68.4$\pm$0.7
Efficiency Analysis. We compare the efficiency of Jump-teaching with the above representative methods and follow the same settings as the synthetic noisy benchmark. The efficiency of these methods is evaluated by throughput and peak memory usage . Throughput is the measurement of the rate at which a method processes picture frames, expressed in thousands of frames per second (kfps). Peak memory usage refers to the maximum amount of memory consumed by a method during its execution. We observe the efficient metrics of these methods with symmetric noise ratio $\epsilon=0.5$ in Table 2, while the accuracy of these methods is illustrated in Table 1. As shown in Table 1 and Table 2, our method achieves almost up to 2.5×2.5$\times$2.5 × speedup, 0.46×0.46$\times$0.46 × peak memory footprint, and superior robustness over all methods. Thus, Jump-teaching achieves optimal empirical results in terms of processing speed, memory usage, and accuracy.
Table 2: Test computational and storage efficiencies.
Standard Decoupling PENCIL Co-teaching Co-teaching+ Topofilter FINE SPRL Jump-teaching
Dual network $\times$ $\times$ $\times$ $\times$ $\times$ $\times$
Throughput (kfps) $\uparrow$ 4.16 1.72 3.13 1.81 2.63 1.61 2.17 2.94 4.07
Peak mem (GB) $\downarrow$ 0.68 1.50 0.73 1.53 1.53 1.40 0.94 0.77 0.70
Real-World Noisy Benchmark. We operate experiments with three trials and report mean value on the Clothing1M dataset. Our proposed method is compared with the following representative approaches: APL, CDR, MentorNet, Decoupling, Co-teaching, Co-teaching+, JoCor, and CoDis. Jump-teaching employs ResNet-50 as the backbone, which is pre-trained on ImageNet[8] dataset and follows the same training setup in [42]. As shown in Table 3, the robustness of Jump-teaching for real-world noisy labels is favorable when compared to other methods.
Table 3: Test accuracy(%) on Clothing1M.
Method APL CDR MentorNet SIGUA Co-teaching Decoupling Co-teaching+ JoCor CoDis Jump-teaching
Accuracy 54.46 66.59 67.25 65.37 67.94 67.65 63.83 69.06 71.60 71.93
4.2 Experiments on Jump-update Strategy
The jump-update strategy is not only suitable for Jump-teaching, but also easily integrated into other methods in LNL. Co-teaching[12] and DivideMix[21] can be regarded as the candidates in supervised-only and semi-supervised methodologies of LNL. We collaborate the jump-update strategy with these two methods respectively.
Collaboration with Supervised-Only Approaches. We utilize ResNet-18 as our backbone architecture. We use only one network from co-teaching and replace cross-update strategy with jump-update strategy, termed J-co-teaching. As Table 4 shows, J-Co-teaching achieves comparable results to Co-teaching with all noise conditions and significantly outperforms Co-teaching in extreme noise scenarios such as $\epsilon=0.8$ and $\epsilon=0.9$.
Table 4: Comparison of test accuracies(%) for Co-teaching and J-Co-teaching on CIFAR-10 and CIFAR-100 with different noise types and ratios.
Dataset CIFAR-10 CIFAR-100
Noise type Sym. Asym. Sym. Asym.
Methods/Noise ratio 0.2 0.5 0.8 0.9 0.4 0.2 0.5 0.8 0.9 0.4
Co-teaching (2*ResNet18) 91.23 87.94 47.69 18.40 85.24 66.22 56.25 20.08 3.82 43.91
J-Co-teaching (ResNet18) 91.52 87.82 74.24 38.41 85.99 66.31 55.39 26.55 9.16 43.53
Collaboration with Semi-Supervised Approaches. DivideMix not only uses two networks to exchange error flow but also employs them to collaboratively predict pseudo-labels. Consequently, we established three baselines in accordance with [21]: DivideMix, DivideMix with $\theta_{1}$ test, and DivideMix w/o co-training. DivideMix with $\theta_{1}$ test utilizes a single network for sample selection, while DivideMix w/o co-training retains a single sample and employs self-update for updating. In each case, we replaced the original update strategy with the jump-update strategy. As Tabel 5 shows, the jump-update strategy significantly improved the accuracy of all baselines, particularly under extreme noise ratios. This indicates that it can effectively overcome accumulated errors without incurring additional costs.
Table 5: Comparison of test accuracies(%) for DivideMix and J-DivideMix on CIFAR-10 and CIFAR-100 with different noise. Best and the mean value of the last ten epochs of accuracy is reported.
Dataset CIFAR-10 CIFAR-100
Noise type Sym. Asym. Sym.
Methods/Noise ratio 0.2 0.5 0.8 0.9 0.4 0.2 0.5 0.8 0.9
DivideMix Best 96.1 94.6 93.2 76.0 93.4 77.3 74.6 60.2 31.5
Last 95.7 94.4 92.9 75.4 92.1 76.9 74.2 59.6 31.0
J-DivideMix Best 96.3 94.9 93.5 78.5 93.0 77.7 74.7 60.0 32.3
Last 96.0 94.6 93.2 77.7 91.9 77.3 74.2 59.7 32.5
DivideMix with $\theta_{1}$ test Best 95.2 94.2 93.0 75.5 92.7 75.2 72.8 58.3 29.9
Last 95.0 93.7 92.4 74.2 91.4 74.8 72.1 57.6 29.2
J-DivideMix with $\theta_{1}$ test Best 96.2 95.0 93.0 80.7 93.3 77.7 73.6 58.5 31.4
Last 95.8 94.8 92.7 79.4 92.2 77.3 73.2 57.8 30.8
DivideMix w/o co-training Best 95.0 94.0 92.6 74.3 91.9 74.8 72.3 56.7 27.7
Last 94.8 93.3 92.2 73.2 90.6 74.1 71.7 56.3 27.2
J-DivideMix w/o co-training Best 95.2 94.5 93.0 80.7 92.4 75.0 72.4 56.3 27.8
Last 94.7 94.0 92.3 79.3 91.3 74.2 71.7 55.3 27.6
5 Conclusion
We propose Jump-teaching, an ultra-efficient and robust algorithm to combat label noise. Our method effectively reduce the accumulated error due to sample-selection bias in a single network which enables our approach to demonstrate outstanding performance across a broad array of experiments, particularly in scenarios characterized by extreme noise conditions. Moreover, with decomposed semantic information of sample losses for selection, our algorithm overcomes the limitations of small-loss criterion and achieve more precise selection. Extensive experimental validation confirms that our method surpasses current state-of-the-art techniques in both accuracy and efficiency. In the future, we hope to expolre whether the network’s selection capability is related to fluctuations in parameters. We did not employ popular semi-supervised techniques to further exploit the identified noisy samples, however, we believe that our method can be further extended by these techniques.
Appendix A Appendix / supplemental material
A.1 Proof of Property 1
The key of proving $D_{A}\propto n$ is to prove $N_{A}=N_{iteration}$. The situation can be categorized into two types: 1) errors accumulate in a single network; 2) errors accumulate across multiple networks.
When errors accumulate in a single network, one accumulation costs one sample selection and one model update. Hence, the single network undergoes one sample selection and parameter update in each iteration.
When errors accumulate across multiple networks, errors first pass through each network once before being transmitted back to the original network. We assume the number of networks denotes $N_{net}$. Error accumulation can only occur when the network’s updates are guided by its own selections. Thus, this process requires $N_{net}$ sample selections and $N_{net}$ parameter updates. Similarly, in each iteration, all networks collectively undergo $N_{net}$ sample selections and $N_{net}$ parameter updates.
Through the analysis of the above two types, Property 1 is proved.
A.2 Discussion on Error Flow
First of all, let’s review the formulation $N_{a}=\frac{N_{A}}{N_{f}}$, mentioned in Property 2. The following discussion is helped by this formulation. We donate the number of samples as $n_{s}$ and the number of mini-batches as $n_{b}$. If the network updates the identifier table after updating the model when the data for selection is a single sample, $N_{f}=n_{s}$. When it comes to mini-batches, $N_{f}=n_{b}$. In all of these conditions, $N_{a}=e$. Thus the single network can be enhanced by integrating the jump-update strategy. If the network updates the identifier table before updating parameters, $N_{f}$ will be doubled so that $N_{a}=0.5e$. The cross-Update strategy uses two networks to mitigate bias, it changed $N_{f}$ to 2. This shows that the cross-update strategy can be incorporated into our framework and does not require the use of two networks as well as the maintenance of two training processes. What’s more, to further reduce $N_{a}$, we provide two methods: the most direct method is to control the update frequency of the identifier table, or by adding new identifier tables to maintain the updates per iteration, each doubling of the identifier table reduces $N_{a}$ by half. This allows us to control $N_{a}$ flexibly. Empirical evidence has shown that maintaining $N_{a}$ in the hundreds is reasonable; too few can lead to the model updating too slowly and over-fitting noise.
A.3 Disccusion on Intial Bias
Initial selection bias and number of accumulations, i.e., $N_{a}$, will jointly decide the overall degree of accumulated error $D_{A}$. Initial Bias is manifested in the form of initial identifier tables which can be obtained in three ways: the first is randomly selection, the second involves updates during warm-up, and the third approach involves labeling before training, which is the usual practice. For ease of subsequent discussion, we call these Random initiation, Warm-up initiation, and Instant initiation. To analyze the impact of initial bias on error accumulation, we observe the subsequent performance of the network by controlling the initial bias, using a single network as the baseline. The update frequency of identifier table i.e., $r$ is set at 10%, and three different initiations are employed. The first two tables are unbiased as they are not affected by selection, whereas Instant initiation is biased as it acquires tables during the second epoch of training and is affected by an increasingly deepening bias. Experimental results are shown in Fig. 4. Identifier table from Random initiation is less accurate than that obtained through Instant initiation, yet its performance is significantly better, demonstrating the effectiveness of avoiding bias. Warm-up initiation gets the best result, as it avoids performance degradation caused by the high error rate of initial table.
Figure 4: Effect of initial bias on training
A.4 The Summary of the Benchmark Datasets
The summary of three benchmark datasets, CIFAR-10, CIFAR-100 and Clothing1M is illustrated in Table 6. Synthetic noisy benchmark in Sec. 4.1 is evaluated on CIFAR-10, and CIFAR-100, respectively, while real-world noisy benchmark in Sec. 4.1 is evaluated on Clothing1M. Specifically, the CIFAR-10 dataset is a standard dataset widely used for image classification tasks. It consists of $60,000$ color images with a resolution of $32\times 32$ pixels, divided into 10 categories, each containing $6,000$ images. The dataset is divided into $50,000$ training images and $10,000$ testing images. Similar to the former, the CIFAR-100 dataset contains more categories. It consists of $60,000$ color images of $32\times 32$ pixels, divided into $100$ categories, with each category containing $600$ images. Clothing1M dataset is a large-scale image classification dataset specifically designed for clothing recognition. It contains over $1,000,000$ labeled images of clothing, covering $14$ different categories such as skirts, shirts, coats, shoes, etc. Images in the Clothing1M dataset typically do not have a fixed standard resolution, as they are sourced from the internet.
Table 6: Summary of datasets used in the experiments.
Dataset # of training # of testing # of class Image size
CIFAR-10 50,000 10,000 10 32x32
CIFAR-100 50,000 10,000 100 32x32
Clothing1M 120,000 10,000 14 not fixed
A.5 Data Augmentation
The network struggles to learn sufficient knowledge from a small number of clean samples with severe noise. Therefore, we adopt a data augmentation strategy following[7] and linearly interpolate the augmented image with its weak view generated by random cropping and horizontal flipping. This enables the network to more easily focus on the key features of the samples. We use Eq. 8 to combine weak and strong views.
$\qquad\boldsymbol{x}_{aug}=\lambda\boldsymbol{x}_{weak}+(1-\lambda)\boldsymbol{x}_{strong}\text{, }\text{ where }\lambda\sim\operatorname{Beta}(\alpha,\alpha).$ (8)
A.6 The Simulation of Noise
As the datasets are clean, we follow [21] and use simulate symmetric noise and asymmetric noise.
Symmetric noise assumes that label errors occur randomly and are independent of the true labels. For a classification problem with $C$ classes, symmetric noise is typically defined as in Eq. 9. The formula includes two things: 1) Each correct label remains unchanged with probability $1-\epsilon+\frac{\epsilon}{C}$. 2) Each correct label $i$ changes to one of the all $C$ classes $j$ with probability $\epsilon$. If the noise is uniformly distributed, the probability of changing to any other class is $\frac{\epsilon}{C}$.
$\begin{cases}1-\epsilon+\frac{\epsilon}{C}&\text{ if }i=j\\ \frac{\epsilon}{C}&\text{ if }i\neq j\end{cases}$ (9)
Asymmetric noise reflects that label errors are more likely to occur between certain classes, typically based on practical scenarios where some classes are visually similar or commonly confused. The probability model for asymmetric noise typically depends on the specific task and data. A simple model might be that the label for a specific class $i$ is incorrectly marked as class $j$ with probability $\epsilon_{ij}$. The asymmetric noise can be formulated in Eq. 10 as follows.
$P(\tilde{Y}=j\mid Y=i)=\epsilon_{ij},$ (10)
where $\epsilon_{ij}$ can be non-zero when $i\neq j$, reflecting specific mislabeling scenarios. where $Y$ represents the true label, $\tilde{Y}$ represents the potentially noisy label, and $\epsilon$ is the noise ratio.
Overall, the symmetric noise assumes that the probability of mislabeling is uniform across all categories, whereas the asymmetric noise considers the similarity between categories. The choice of noise typically depends on the characteristics of the dataset and the specific requirements of the task.
A.7 A Brief Overview of Compared Methods
We reimplemented Decoupling, Co-teaching, Co-teaching+, and PENCIL using the publicly available code***https://github.com/JackYFL/DISC. Similarly, FINE, SPRL, and TopoFilter were re-implemented using their respective official repositorieshttps://github.com/Kthyeon/FINE_officialhttps://github.com/pxiangwu/TopoFilter §§§https://github.com/xsshi2015/Self-paced-Resistance-Learning. Results presented in Table 3 are cited from [42]. The outcomes for Decoupling, Co-teaching+, and PENCIL on CIFAR-10 and CIFAR-100 with symmetric noise, shown in Table 1, are derived from [22]. Additionally, baseline results featured in 5 for DivideMix, DivideMix with $\theta_{1}$ test, and DivideMix without co-training are sourced from [21].
A.8 The Structures of Backbones and the Auxiliary Head
The auxiliary head of Jump-teaching consists of three fully connected layers with ReLU activations and dropout layers in between, ending with a Tanh activation. For CIFAR-10 and Clothing1M, the output dimension of the last fully connected layer is set to $32$, whereas setting to $64$ for CIFAR-100. The intermediate dimension is uniformly set to $512$.
A.9 Analysis of Threshold
We test the sensitivity of Jump-teaching to the threshold. The settings maintain the same and we conduct experiments on CIFAR-10 with symmetric noise ratio $\epsilon=0.5$ and asymmetric noise ratio $\epsilon=0.4$. The results in Table 7 indicate a relatively stable performance across different thresholds, demonstrating that the Jump-teaching method’s effectiveness is not significantly influenced by the exact value of $\tau$ and is more adaptable to various datasets. We attribute the success to two main factors. First, the Hadamard matrix used for constructing hashing codes exhibits uniformness in each column[45], a feature that ensures variance consistently reflects the required patterns. Second, we utilize classification heads to collaboratively select samples, ensuring a sufficient number of samples are always available for learning. These combined elements significantly enhance the stability and accuracy of the model across different datasets.
Table 7: Test Accuracies(%) with different threshold settings
$\tau$ 0.001 0.005 0.01 0.05 0.1
Sym. $\epsilon=0.5$ 92.12 92.17 92.04 91.97 92.17
Asym. $\epsilon=0.4$ 90.71 90.66 90.79 90.82 90.70
A.10 Ablation Study
We evaluate three components of Jump-teaching: data augmentation(AUG), jump-update strategy(JU) and sample selection(SS). Results of ablation study are shown in Table 8. Effectiveness of Data Augmentation When only a limited number of samples are available, such as in the presence of symmetric noise ratio $\epsilon=0.8$, data augmentation can significantly enhance the learning capabilities of these samples, enabling the network to focus more on the intrinsic characteristics of the images themselves. Effectiveness of Jump-update Strategy The jump-update strategy can effectively overcome error accumulation, allowing it to perform excellently with extreme noise. Without the jump-update strategy, the network is almost unable to learn with the symmetric noise ratio $\epsilon=0.8$ condition on CIFAR-10.
Effectiveness of Sample Selection If noisy samples are not filtered out at all, the network can easily fit them, resulting in reduced generalization performance. This has been verified with various noise conditions.
Table 8: Ablation Study
Modules CIFAR-10 CIFAR-100
AUG JU SS Sym. 20% Sym. 50% Sym. 80% Asym. 40% Sym. 20% Sym. 50% Sym. 80% Asym. 40%
$\times$ $\checkmark$ $\checkmark$ 94.12 89.59 74.94 85.34 74.69 65.87 32.66 67.30
$\checkmark$ $\times$ $\checkmark$ 88.33 81.27 10.01 56.28 68.32 65.27 28.71 50.40
$\checkmark$ $\checkmark$ $\times$ 85.42 63.21 30.18 78.19 61.88 39.35 12.90 45.17
$\times$ $\times$ $\checkmark$ 92.21 87.24 9.99 57.56 74.19 65.98 29.23 57.05
$\checkmark$ $\checkmark$ $\checkmark$ 94.77 92.11 82.81 90.91 72.15 66.91 40.81 67.66
A.11 Discussion about Disaggrement
We conduct experiments on CIFAR-10 symmetric noise ratio $\epsilon=0.5$, $\epsilon=0.8$, using a three-layer MLP. The method of selecting clean data follows Co-teaching. We use Intersection over Union (IoU) between two selections to represent the disagreement. Specifically, Intersection computes the number of common selected samples between two, while Union calculates the total number of selected samples between two. In the cross-update strategy, the disagreement emerges in a different network, while in the other two strategies, it emerges during different epochs. As can be seen from Fig. 1(1(c)) and Fig. 1(1(d)), Jump-Update and Self-Update The single network exhibits significantly greater disagreement between epochs than the dual-network setup, which we aim to utilize effectively. The Jump-Update, by leveraging previous selections to update the model, shows slightly greater disagreement compared to self-update. Simultaneously, we observe that the error band of Jump-Update is smaller than the other two methods, indicating more stable network updates. This reflects the network’s robustness against noise to some extent. These results are consistent with Fig. 1(1(a)) and Fig. 1(1(b)): Jump-Update, with the smallest error band, achieves the best performance, while self-update, with the largest error band, performs the worst.
A.12 The Choice of Jump Step.
We test the effect of different jump steps on Jump-teaching with the results shown in Table 9. Performance across various steps remains similar as error accumulation has been effectively constrained to the epoch level. In the Jump-update Strategy, we opt for a step size of 2, obviating the need for additional forward propagation post-training. Furthermore, empirical evidence indicates a marginal decline in performance as step size increases, corroborating the assertions in Sec. 2 that delays in updating tables can detrimentally impact the network’s capacity to assimilate newly introduced clean samples.
Table 9: Test accuracy(%) on CIFAR-10 and CIFAR-100 with symmetric and asymmetric noise with various noise settings
Noise type and ratio/ Step 1 2 3 4 5 10 15 20 25 30
CIFAR-10 Sym. $\epsilon=0.8$ 82.81 83.37 82.47 83.09 82.55 82.04 81.03 81.35 81.46 79.77
CIFAR-10 Asym. $\epsilon=0.4$ 90.91 90.77 90.66 90.78 90.86 90.42 90.35 90.60 90.69 90.23
CIFAR-100 Sym. $\epsilon=0.5$ 66.91 67.03 66.85 66.21 66.27 65.54 65.87 65.88 65.46 64.67
A.13 The Size of Hadamard Codebook.
As demonstrated in Table 10, we find that the network’s performance does not exhibit a clear correlation with the code length. This indicates that our approach differs from traditional deep hashing methods[45, 48, 18] which rely on learning semantic features of samples. Typically, the performance of such methods is closely associated with the dimensionality of the high-dimensional space, as more dimensions typically provide richer information. In contrast, our method employs non-mutually exclusive encoding to reflect patterns of memorization effects, showcasing the robustness of our approach.
Table 10: Test accuracy(%) on CIFAR-10 and CIFAR-100 with symmetric and asymmetric noise.
Dataset CIFAR-10 CIFAR-100
Noise type Sym. Asym. Sym. Asym.
Code length/Noise ratio 0.2 0.5 0.8 0.4 0.2 0.5 0.8 0.4
8 94.66 91.85 82.75 90.56 - - - -
16 94.7 92.00 82.87 90.79 - - - -
32 94.74 92.3 83.28 90.99 - - - -
64 94.67 92.10 83.41 90.51 72.14 66.91 40.81 67.66
128 94.62 91.95 83.1 90.55 72.04 66.39 38.25 68.86

### v4

@@
 In the cross-update strategy, the disagreement emerges in a different network, while in the other two strategies, it emerges during different epochs.
-More details are discussed in Appendix A.11
+More details are discussed in Appendix A.15.
 Sample selection can be regarded as a potential method for LNL[34, 19, 40, 28, 12, 42].
@@
 Our contributions can be summarized as follows: 1) We exploit the disagreement between intervalic iterations during training for a single network, and leverage the jump-update strategy to suppress the accumulated error.
-To the best of our knowledge, it is the first time that a single network has surpassed the capacity of a dual network in the robustness and efficiency of LNL.
+To the best of our knowledge, it is the first time that a single network outperforms a dual network in terms of robustness and efficiency in LNL.
 (2) The loss of sample selection is proposed with a lightweight design and easily selects clean data by intrinsic characteristics.
@@
 Furthermore, we offer both theoretical and experimental analysis to explain its effectiveness and then explore its relationship with other approaches.
-Strategy Description sTo simplify our discussion, we dive into the procedure of sample selection and give some definitions.
+Strategy Description.
+To simplify our discussion, we dive into the procedure of sample selection and give some definitions.
 Sample selection is accomplished by the execution of model updates and clean sample selection with multiple iterations.
@@
 4.2.
-Theoretical Analysis Sample selection inevitably has a selection bias, leading to accumulated error in an error flow.
+Empirical Analysis.
+Sample selection inevitably has a selection bias, leading to accumulated error in an error flow.
 As shown in Fig.
@@
 This is discussed in Appendix A.4, which also includes the impact of initial bias.
-Conclusion The jump-update strategy is more effective and efficient than previous works.
+Conclusion.
+The jump-update strategy is more effective and efficient than previous works.
 Compared to methods that implicitly leverage Property 1, it can reduce $D_{a}$ to a small number i.e., $0.5e$ at a constant cost.
@@
 3.2 Semantic Loss Decomposition
-Motivation According to memorization effects, neural networks prioritize learning simpler patterns from data [49].
+Motivation.
+According to memorization effects, neural networks prioritize learning simpler patterns from data [49].
 The small-loss criterion leverages this effect: samples with clean labels are learned first by the network, hence exhibiting smaller losses compared to noisy samples.
@@
 The notation $\mathbf{0}$ denotes a vector of zeros, indicating that the distances across different bits uniformly converge towards zero as $t$ approaches infinity.
-Hash Coding with Hardamard Matrix Inspired by [45], we utilize the excellent properties of Hadamard matrices to construct mappings for category encoding.
+Hash Coding with Hardamard Matrix.
+Inspired by [45], we utilize the excellent properties of Hadamard matrices to construct mappings for category encoding.
 A $K$-bit Hadamard matrix can generate $2K$ codewords, each $K$ bits long, with a minimum Hamming distance of $\frac{K}{2}$.
@@
 This mapping can be presented as $H:\tilde{y}_{i}\in\{0,1\}^{C}\rightarrow\tilde{y}_{i}^{\prime}\in\{-1,1\}^{k}$.
-Feature Embedding and Loss Function outputs are mapped to feature embeddings in a subspace via an additional three-layer MLP as detection head, which represented as $f:x_{i}\in\mathbb{R}^{n}\rightarrow z_{i}\in\mathbb{R}^{k}$.
+Feature Embedding and Loss Function.
+Outputs are mapped to feature embeddings in a subspace via an additional three-layer MLP as detection head, which represented as $f:x_{i}\in\mathbb{R}^{n}\rightarrow z_{i}\in\mathbb{R}^{k}$.
 In this framework, we employ binary cross-entropy[31] to define the distance $\mathbf{d}$ between the network’s predictions and the labels, which indicates whether the network adequately captures the semantics of each component, which is formulated as
@@
 3:     Permute $D$ randomly
-4:     for $s=1$ to $I_{\text{max}}$ do
+4:     for $s=1$ to $S_{\text{max}}$ do
 5:        $clean\_flag\leftarrow I[s]$
@@
 Noise ratio 0.2 0.5 0.8 0.4 0.2 0.5 0.8 0.4
-Standard 84.6 $\pm$ 0.1 62.4 $\pm$ 0.3 27.3$\pm$0.3 75.92 $\pm$0.4 56.1 $\pm$0.1 33.6 $\pm$ 0.2 8.2 $\pm$0.1 40.1 $\pm$0.2
+Standard 84.6$\pm$0.1 62.4$\pm$0.3 27.3$\pm$0.3 75.9$\pm$0.4 56.1$\pm$0.1 33.6$\pm$0.2 8.2$\pm$0.1 40.1$\pm$0.2
 Decoupling(’17) 86.4$\pm$0.1 72.9$\pm$0.2 48.4$\pm$0.6 83.3$\pm$0.2 53.3$\pm$0.1 28.0$\pm$0.1 7.9$\pm$0.1 39.9$\pm$0.4
 Co-teaching(’18) 89.9$\pm$0.6 67.3$\pm$4.2 28.1$\pm$2.0 79.2$\pm$0.5 61.8$\pm$0.4 34.7$\pm$0.5 7.5$\pm$0.5 40.0$\pm$1.2
-Co-teaching+(’19) 88.1 $\pm$ 0.0 61.8 $\pm$ 0.2 22.3 $\pm$ 0.6 58.2$\pm$ 0.2 54.5 $\pm$ 0.1 27.6 $\pm$ 0.1 8.4 $\pm$ 0.1 19.87$\pm$0.3
+Co-teaching+(’19) 88.1$\pm$0.0 61.8$\pm$0.2 22.3$\pm$0.6 58.2$\pm$0.2 54.5$\pm$0.1 27.6$\pm$0.1 8.4$\pm$0.1 19.9$\pm$0.3
 PENCIL(’19) 88.2$\pm$0.6 73.4$\pm$1.5 36.0$\pm$0.7 77.3$\pm$3.4 57.4$\pm$1.0 11.4$\pm$3.2 5.4$\pm$1.2 45.7$\pm$0.9
@@
 FINE(’21) 90.2$\pm$0.1 85.8$\pm$0.8 70.8$\pm$1.8 87.8$\pm$0.1 70.1$\pm$0.3 57.9$\pm$1.2 22.2$\pm$0.7 53.5$\pm$0.8
-SPRL(’23) 91.7 $\pm$ 0.2 88.35 $\pm$ 0.6 63.9 $\pm$ 1.4 89.8 $\pm$ 0.5 69.5 $\pm$ 0.8 57.17 $\pm$ 1.6 23.8 $\pm$ 2.2 58.72 $\pm$ 1.3
+SPRL(’23) 91.7$\pm$0.2 88.4$\pm$0.6 63.9$\pm$1.4 89.8$\pm$0.5 69.5$\pm$0.8 57.2$\pm$1.6 23.8$\pm$2.2 58.7$\pm$1.3
 Jump-teaching 94.8$\pm$0.1 92.2$\pm$0.1 84.1$\pm$1.1 90.7$\pm$0.3 72.7$\pm$0.5 67.1$\pm$0.2 40.0$\pm$1.1 68.4$\pm$0.7
@@
 Methods/Noise ratio 0.2 0.5 0.8 0.9 0.4 0.2 0.5 0.8 0.9 0.4
+Co-teaching (1*ResNet18) 88.55 84.39 43.21 15.88 78.27 61.61 50.97 16.58 3.61 39.63
 Co-teaching (2*ResNet18) 91.23 87.94 47.69 18.40 85.24 66.22 56.25 20.08 3.82 43.91
-J-Co-teaching (ResNet18) 91.52 87.82 74.24 38.41 85.99 66.31 55.39 26.55 9.16 43.53
+J-Co-teaching ( 1*ResNet18) 91.52 87.82 74.24 38.41 85.99 66.31 55.39 26.55 9.16 43.53
 Collaboration with Semi-Supervised Approaches.
@@
 Appendix A Appendix / supplemental material
-A.1 Proof of Property 1
+A.1 Workflow of Jump-teaching
+Figure 4: The framework and workflow of Jump-teaching.
+Motivated by two key issues in the sample selection method, Jump-teaching proposes a novel model-update strategy to mitigate the selection bias and a plug-and-play sample selection component to replace complex modeling.
+A.2 Proof of Property 1
 The key of proving $D_{A}\propto n$ is to prove $N_{A}=N_{iteration}$.
@@
 Through the analysis of the above two types, Property 1 is proved.
-A.2 Discussion on Error Flow
+A.3 Discussion on Error Flow
 First of all, let’s review the formulation $N_{a}=\frac{N_{A}}{N_{f}}$, mentioned in Property 2.
@@
 Empirical evidence has shown that maintaining $N_{a}$ in the hundreds is reasonable; too few can lead to the model updating too slowly and over-fitting noise.
-A.3 Disccusion on Intial Bias
+A.4 Disccusion on Intial Bias
 Initial selection bias and number of accumulations, i.e., $N_{a}$, will jointly decide the overall degree of accumulated error $D_{A}$.
@@
 Experimental results are shown in Fig.
-4.
+5.
 Identifier table from Random initiation is less accurate than that obtained through Instant initiation, yet its performance is significantly better, demonstrating the effectiveness of avoiding bias.
 Warm-up initiation gets the best result, as it avoids performance degradation caused by the high error rate of initial table.
-Figure 4: Effect of initial bias on training
-A.4 The Summary of the Benchmark Datasets
+Figure 5: Effect of initial bias on training.
+A.5 Loss Convergence
+We present results for four different temperature values to show how the different values of Temperature scaling affect the convergence of the losses.
+When T=1, the Binary Cross-Entropy (BCE) loss fails to converge properly.
+The cases for T=2, 3 and 4 illustrate the effects of temperature scaling.
+Specifically, temperature scaling slows down the convergence speed of the classification head, which can be obviously observed during the warm-up phase, while it accelerates convergence of the detection head which can be obviously observed during the training phase.
+Notably, as the temperature T increases, the convergence rate of the Cross-Entropy (CE) loss decreases.
+(a) Convergence of Loss BCE
+(b) Convergence of Loss CE
+Figure 6: Convergence of losses with different temperature scaling factors $T$ under CIFAR-100 Sym. $\epsilon=0.8$.
+A.6 Details of Sample Selecion
+We include the number and ratio of clean samples selected in each epoch.
+This data is reported for each noise condition in the CIFAR-10 and CIFAR-100 datasets and is presented in Fig.
+7.
+It can be observed that the number of clean samples continuously increases and stabilizes during the last 20 epochs, presented in Fig.
+7(a) and Fig.
+7(c).
+As for Fig.
+7(b) and Fig.
+7(d), the ratio of clean samples increases steadily during the early and middle stages of training, indicating that the network’s selection ability improves as its performance increases.
+However, in the later stages (epochs range from 150 to 200), the ratio of clean samples begins to decline, primarily due to the model starting to fit the noisy data, which results in more noisy samples being misclassified as clean.
+(a) Clean num. on CIFAR-10
+(b) Clean rat. on CIFAR-10
+(c) Clean num. on CIFAR-100
+(d) Clean rat. on CIFAR-100
+Figure 7: The clean number and the clean ratio of selected samples on CIFAR-10 and CIFAR-100 with different noise types and ratios.
+A.7 More Experiments on Different Noise Types
+We have added the experiments of pairflip-45 and instance-dependent label noise i.e.
+IDN.
+Results are shown in Tab.
+6 and Tab.
+7.
+Coupled with the noises in the original paper, our experiments support up to five types of noise.
+On pairflip noise, our algorithm also exhibits its strong performance, with an increase of 6.73% and 9.95% on CIFAR-10 and CIFAR-100, respectively.
+Our algorithm maintains a huge advantage in instance-dependent noise.
+The leading margins are 3.95% (CIFAR-10 IDN.
+0.2), 6.74% (CIFAR-10 IDN.
+0.4), 10.2% (CIFAR-10 IDN.
+0.6), 4.87% (CIFAR-100 IDN.
+0.2), 4.3% (CIFAR-100 IDN.
+0.4) and 12.67% (CIFAR-100 IDN.
+0.6).
+This indicates that our algorithm is also extremely robust under IDN settings.
+Table 6: Performance of methods on CIFAR-10 and CIFAR-100 with pairflip $\epsilon=0.45$ noise.
+Methods CIFAR-10 CIFAR-100
+CE 50.22 $\pm$ 0.43 21.59 $\pm$ 0.87
+Co-teaching 52.99 $\pm$ 0.16 33.22 $\pm$ 0.66
+Co-teaching+ 55.19 $\pm$ 0.27 29.26 $\pm$ 0.15
+PENCIL 54.58 $\pm$ 2.26 24.97 $\pm$ 0.57
+SPRL 90.54 $\pm$ 0.02 53.62 $\pm$ 1.07
+FINE 77.09 $\pm$ 0.10 41.62 $\pm$ 1.01
+Topo 84.60 $\pm$ 0.45 52.40 $\pm$ 1.42
+Jump-teaching 91.33 $\pm$ 0.23 62.35 $\pm$ 0.74
+Table 7: Performance of methods on CIFAR-10 and CIFAR-100 with varying IDN noise ratios.
+Methods CIFAR-10 CIFAR-100
+0.2 0.4 0.6 0.2 0.4 0.6
+CE 85.45 $\pm$ 0.57 76.23 $\pm$ 1.54 59.75 $\pm$ 1.30 57.79 $\pm$ 1.25 41.15 $\pm$ 0.83 25.68 $\pm$ 1.55
+Co-teaching 88.87 $\pm$ 0.24 73.00 $\pm$ 1.24 62.51 $\pm$ 1.98 43.30 $\pm$ 0.39 23.21 $\pm$ 0.57 12.58 $\pm$ 0.51
+Co-teaching+ 89.80 $\pm$ 0.28 73.78 $\pm$ 1.39 59.22 $\pm$ 6.34 41.71 $\pm$ 0.78 24.45 $\pm$ 0.71 12.58 $\pm$ 0.51
+PENCIL 86.23 $\pm$ 1.23 69.88 $\pm$ 1.39 44.16 $\pm$ 1.58 57.41 $\pm$ 0.69 48.38 $\pm$ 0.52 29.69 $\pm$ 3.03
+FINE 90.85 $\pm$ 1.21 84.53 $\pm$ 0.77 61.24 $\pm$ 2.02 60.14 $\pm$ 0.015 43.05 $\pm$ 1.43 28.88 $\pm$ 2.01
+Topo 89.66 $\pm$ 0.86 85.95 $\pm$ 1.02 75.28 $\pm$ 0.64 64.41 $\pm$ 0.37 55.00 $\pm$ 0.82 31.61 $\pm$ 1.45
+SPRL 87.99 $\pm$ 0.03 86.15 $\pm$ 0.81 75.01 $\pm$ 0.74 67.66 $\pm$ 0.17 64.55 $\pm$ 0.19 47.16 $\pm$ 1.97
+Jump-teaching 94.80 $\pm$ 0.28 92.89 $\pm$ 0.33 85.48 $\pm$ 0.41 72.53 $\pm$ 0.21 68.85 $\pm$ 0.27 59.83 $\pm$ 0.09
+A.8 The Summary of the Benchmark Datasets
 The summary of three benchmark datasets, CIFAR-10, CIFAR-100 and Clothing1M is illustrated in Table 8.
@@
 Clothing1M 120,000 10,000 14 not fixed
-A.5 Data Augmentation
+A.9 Data Augmentation
 The network struggles to learn sufficient knowledge from a small number of clean samples with severe noise.
@@
 $\qquad\boldsymbol{x}_{aug}=\lambda\boldsymbol{x}_{weak}+(1-\lambda)\boldsymbol{x}_{strong}\text{, }\text{ where }\lambda\sim\operatorname{Beta}(\alpha,\alpha).$ (8)
-A.6 The Simulation of Noise
+A.10 The Simulation of Noise
 As the datasets are clean, we follow [21] and use simulate symmetric noise and asymmetric noise.
@@
 The choice of noise typically depends on the characteristics of the dataset and the specific requirements of the task.
-A.7 A Brief Overview of Compared Methods
+A.11 A Brief Overview of Compared Methods
 We reimplemented Decoupling, Co-teaching, Co-teaching+, and PENCIL using the publicly available code***https://github.com/JackYFL/DISC.
@@
 Additionally, baseline results featured in 5 for DivideMix, DivideMix with $\theta_{1}$ test, and DivideMix without co-training are sourced from [21].
-A.8 The Structures of Backbones and the Auxiliary Head
+A.12 The Structures of Backbones and the Auxiliary Head
 The auxiliary head of Jump-teaching consists of three fully connected layers with ReLU activations and dropout layers in between, ending with a Tanh activation.
@@
 The intermediate dimension is uniformly set to $512$.
-A.9 Analysis of Threshold
+A.13 Analysis of Threshold
 We test the sensitivity of Jump-teaching to the threshold.
@@
 These combined elements significantly enhance the stability and accuracy of the model across different datasets.
-Table 7: Test Accuracies(%) with different threshold settings
+Table 9: Test Accuracies(%) with different threshold settings.
 $\tau$ 0.001 0.005 0.01 0.05 0.1
@@
 Asym. $\epsilon=0.4$ 90.71 90.66 90.79 90.82 90.70
-A.10 Ablation Study
-We evaluate three components of Jump-teaching: data augmentation(AUG), jump-update strategy(JU) and sample selection(SS).
-Results of ablation study are shown in Table 8.
-Effectiveness of Data Augmentation When only a limited number of samples are available, such as in the presence of symmetric noise ratio $\epsilon=0.8$, data augmentation can significantly enhance the learning capabilities of these samples, enabling the network to focus more on the intrinsic characteristics of the images themselves.
-Effectiveness of Jump-update Strategy The jump-update strategy can effectively overcome error accumulation, allowing it to perform excellently with extreme noise.
+A.14 Ablation Study
+We evaluate three components of Jump-teaching: data augmentation(AUG), jump-update strategy(JU) and sample selection(SS). Results of ablation study are shown in Table 10.
+Effectiveness of Data Augmentation.
+When only a limited number of samples are available, such as in the presence of symmetric noise ratio $\epsilon=0.8$, data augmentation can significantly enhance the learning capabilities of these samples, enabling the network to focus more on the intrinsic characteristics of the images themselves.
+Effectiveness of Jump-update Strategy.
+Jump-update strategy can effectively overcome error accumulation, allowing it to perform excellently with extreme noise.
 Without the jump-update strategy, the network is almost unable to learn with the symmetric noise ratio $\epsilon=0.8$ condition on CIFAR-10.
-Effectiveness of Sample Selection If noisy samples are not filtered out at all, the network can easily fit them, resulting in reduced generalization performance.
-This has been verified with various noise conditions.
-Table 8: Ablation Study
+Effectiveness of Sample Selection.
+If noisy samples are not filtered out at all, the network can easily fit them, resulting in reduced generalization performance.
+This has been verified with various noises.
+Table 10: Ablation Study.
 Modules CIFAR-10 CIFAR-100
@@
 $\checkmark$ $\checkmark$ $\checkmark$ 94.77 92.11 82.81 90.91 72.15 66.91 40.81 67.66
-A.11 Discussion about Disaggrement
+A.15 Discussion about Disaggrement
 We conduct experiments on CIFAR-10 symmetric noise ratio $\epsilon=0.5$, $\epsilon=0.8$, using a three-layer MLP.
@@
 1(1(b)): Jump-Update, with the smallest error band, achieves the best performance, while self-update, with the largest error band, performs the worst.
-A.12 The Choice of Jump Step.
-We test the effect of different jump steps on Jump-teaching with the results shown in Table 9.
+A.16 The Choice of Jump Step
+We test the effect of jump steps on Jump-teaching with the results shown in Table 11.
 Performance across various steps remains similar as error accumulation has been effectively constrained to the epoch level.
@@
 2 that delays in updating tables can detrimentally impact the network’s capacity to assimilate newly introduced clean samples.
-Table 9: Test accuracy(%) on CIFAR-10 and CIFAR-100 with symmetric and asymmetric noise with various noise settings
+Table 11: Test accuracy(%) on CIFAR-10 and CIFAR-100 with symmetric and asymmetric noise with various noise settings.
 Noise type and ratio/ Step 1 2 3 4 5 10 15 20 25 30
@@
 CIFAR-100 Sym. $\epsilon=0.5$ 66.91 67.03 66.85 66.21 66.27 65.54 65.87 65.88 65.46 64.67
-A.13 The Size of Hadamard Codebook.
+A.17 The Size of Hadamard Codebook
 As demonstrated in Table 12, we find that the network’s performance does not exhibit a clear correlation with the code length.